A copper wire has a resistance of 15.0 at a temperature of 14 degrees * C . After a current passes through the windings, the resistance rose to 22.5 ohms. To what temperature was the wire heated?


1
Expert's answer
2022-04-07T18:25:44-0400

"t_2=t_1+\frac{R_2-R_1}{\alpha R_1}=146°C."